# Stolz positive scalar curvature structure groups, proper actions and equivariant 2-types

### v1

@@ Abstract
In this note, we study equivariant versions of Stolz’ R-groups, the positive scalar curvature structure groups $R_{n}(X)^{G}$, for proper actions of discrete groups $G$. We define the concept of a fundamental groupoid functor for a $G$-space, encapsulating all the fundamental group information of all the fixed point sets and their relations. We construct classifying spaces for fundamental groupoid functors. As a geometric result, we show that Stolz’ equivariant R-group $R_{n}(X)^{G}$ depends only on the fundamental groupoid functor of the reference space $X$. The proof covers at the same time in a concise and clear way the classical non-equivariant case.
@@ 1 Introduction
$\ldots\to\mathrm{R}^{spin}_{n+1}(X)\to\mathrm{Pos}^{spin}_{n}(X)\to\Omega^{spin}_{n}(X)\to\mathrm{R}^{spin}_{n}(X)\to\ldots$ (1.1)
of cobordism groups, where $X$ is a CW-complex. This long exact sequence includes the well-known spin cobordism $\Omega_{*}^{spin}(X)$; then the structure bordism group of metrics of positive scalar curvature $\mathrm{Pos}_{*}^{spin}(X)$, whose cycles $(f\colon M\to X,g)$ are defined like those of $\Omega_{*}^{spin}(X)$, but adding the geometric secondary structure given by a positive scalar curvature metric $g$ on the smooth compact manifold $M$ with spin structure; finally, the object of study of this note, $\mathrm{R}^{spin}_{*}(X)$ which arises as the homotopy fiber of the forgetful functor $\mathrm{Pos}^{spin}_{*}(X)\to\Omega^{spin}_{*}(X)$.
An explicit calculation of this last group has not yet been achieved in any case of interest. Nonetheless, Stolz has proved that, if $X=B\pi_{1}(M)$ is the classifying space of the fundamental group of a closed spin manifold $M$, then $\mathrm{R}^{spin}_{\dim M+1}(X)$ acts freely and transitively on the space of concordance classes of positive scalar curvature metrics on $M$, which then inherits canonically the structure of affine space modelled on an abelian group. This has been used, for instance, in [PSZ, PSZ2, SZ, XieYuZeidler] to give a lower bound on the rank of this affine group and of the moduli space of concordance classes of metrics of positive scalar curvature, where moduli space refers to the quotient by the action of the diffeomorphism group of $M$.
A fundamental step in [SZ] uses the fact, which can be implicitly deduced from [Stolz], that a 2-connected map between CW-complexes, such as the classifying map $u\colon M\to B\pi_{1}(M)$, induces an isomorphism between $\mathrm{R}^{spin}_{*}$ groups.
There is an obvious equivariant reformulation of this story: instead of working with a space $X$ one can work with the a $G$-cover $\bar{X}$ with its free $G$-action. One can then go equivalently go back and forth between structure on $X$ and $G$-invariant structure on $\bar{X}$ (pulling back from $X$ to $\bar{X}$ and quotienting by $G$ from $\bar{X}$ to $X$). This way, for a space $X$ with fundamental group $G$, one can replace equivalently the Stolz sequence (1.1) by its equivariant version
@@
where all the cycles are defined as above, by required to come with a free and co-compact and (whenever we have metrics) isometric action of $G$, and all maps are required to be $G$-equivariant. This equivalent reformulation is important to get information about the groups in the Stolz sequence via higher index theory, which has been successfully implemented in [PiazzaSchick] and [Zeidler], compare also [XieYu]. The proofs of lower bounds on the rank of the affine group of concordance classes of positive scalar curvature are based on these techniques.
The aim of this note is two-fold. For once, we want to give an explicit and concise proof of the fact that $R^{spin}_{*}(X)$ does only depend on the fundamental group information of $X$. Secondly, we want to analyze this for the case of general proper $G$-actions. In particular, we want to develop basic tools for this generalization and then study it in the equivariant context of a CW-complex endowed with a proper action of a discrete group.
@@ A $G$-CW-complex X is a $G$-space together with a $G$-invariant filtration
Let $f\colon X\to Y$ be a $G$-map. Then there exists a G-homotopy $h\colon X\times I\to Y$ such that $h_{0}=f$ and $h_{1}$ is cellular.
@@
(compare [ttd, (2.6) Proposition]) Let $f\colon B\to C$ be a $\nu$-connected map between $G$-CW-complexes and $A$ another $G$-CW-complex. Write $[A,B]^{G}$ for the set of $G$-homotopy classes of $G$-maps from $A$ to $B$. Then
@@
As a particular case of this proposition we obtain the following result [ttd, (2.7) Proposition].
Proposition 2.5.
Let $f\colon Y\to Z$ be a $G$-map between $G$-complexes such that $f^{H}$ is a homotopy equivalence for all subgroups $H$ of $G$. Then $f$ is a $G$-homotopy equivalence.
@@ 3 The Stolz exact sequence
  • $R^{spin}_{n}(X)^{G}$ is the bordism group of spin $G$-manifolds with boundary (possibly empty) of dimension $n$, together with a $G$-invariant Riemannian metric of positive scalar curvature on the boundary. Bordisms are then manifolds with corners. In particular, $(M,f,g)$ and $(M^{\prime},f^{\prime},g^{\prime})$ are equivalent if there exists a $G$-bordism $(W,F,\bar{g})$, where $W$ is a bordism between $M$ and $M"$ and the resulting bordism $\partial_{0}W$ between $\partial M$ and $\partial M^{\prime}$, $\bar{g}$ carries a $G$-invariant metric with positive scalar curvature, so that it is a bordism between $(\partial M,\partial f,g)$ and $(\partial M^{\prime},\partial f^{\prime},g^{\prime})$ in the sense of $Pos^{spin}_{n-1}(X)^{G}$.
@@
Note that for each cycle $f\colon M\to X$ for $\Omega_{n}^{spin}(X)^{G}$ the isotropy of $M$ is restricted to belong to the family $\mathcal{I}(X)$ (because the image under the equivariant map $f$ of a point $x\in M$ fixed by a subgroup $H$ of $G$ must also be fixed by $H$, and is a point in $X$).
@@
[…]
@@
This is a rather direct consequence of the definitions. The proof for the $G$-equivariant case is exactly the same as the classical non-equivariant situation. ∎
@@
    It follows that induction give an isomorphism (already on the level of cycles and relations)
@@
The isotropy family $\mathcal{I}(\tilde{X})$ of the action of $\widetilde{G}$ on $\tilde{X}$ is precisely the inverse image of $\mathcal{I}(X)$ under the projection $\widetilde{G}\to G$.
This again follows from covering theory: if we fix $x\in X$ with lift $\tilde{x}\in\tilde{X}$ and a subgroup $H$ of $G$ fixing $x$ we have a canonical split $H\to\widetilde{G}$ sending $h\in H$ to the unique pair $(\alpha\colon\tilde{X}\to\tilde{X},h)$ such that $\alpha(\tilde{x})=\tilde{x}$.
@@
The action of $\pi_{1}(X)\subset\widetilde{G}$ on $\tilde{X}$ is free. Consequently, for every $\widetilde{G}$-map $f\colon\bar{M}\to\tilde{X}$ the action of $\pi_{1}(X)$ on $\bar{M}$ is free. We can therefore quotient out this action and obtain a cycle $M:=\bar{M}/\pi_{1}(X)\to X$ with residual action of $G:=\widetilde{G}/\pi_{1}(X)$. Vice versa, given a $G$-map $f\colon M\to X$ we can pull back $\tilde{X}\to X$ along $f$ and obtain a $\pi_{1}(X)$-covering $\tilde{M}\to M$ with map $\tilde{f}\colon\tilde{M}\to\tilde{X}$. Because $f$ is a $G$-map, it is straightforward to pull back also the action of $\widetilde{G}$ to an action on $\tilde{M}$ which covers the action of $G$ on $M$ and such that $\tilde{f}$ is a $\widetilde{G}$-map (indeed, the $\widetilde{G}$-action on $\tilde{M}$ is defined mapping a point $(m,\tilde{x})\in\tilde{M}\subset M\times\tilde{X}$ by $(\alpha,g)\in\widetilde{G}$ to $(gm,\alpha(\tilde{x}))$).
@@
For the paper at hand, the transition to the universal covering as in Proposition 3.6 is not really relevant. However, for the classical Stolz positive scalar curvature exact sequence, all efficient ways to get information about it (and consequently about sets of positive scalar curvature on a compact manifold $M$) use higher index theory and pass through the equivariant version, as we have already mentioned in the introduction. It is pretty obvious that the same observation does hold for non-free actions: even if we want to understand symmetric metrics of positive scalar curvature on a compact manifold $M$ with smooth action of a finite group $\Gamma$, it will be very useful to pass to the covering space which also sees the symmetries induced by the non-trivial fundamental group. This has already been implemented, and the map from the equivariant Stolz sequence to analysis in the sense of [PiazzaSchick] has been constructed in [XieYu, Section 5] and used in [XieYuZeidler].

### v2

@@ Abstract
In this note, we study equivariant versions of Stolz’ R-groups, the positive scalar curvature structure groups $R^{spin}_{n}(X)^{G}$, for proper actions of discrete groups $G$. We define the concept of a fundamental groupoid functor for a $G$-space, encapsulating all the fundamental group information of all the fixed point sets and their relations. We construct classifying spaces for fundamental groupoid functors. As a geometric result, we show that Stolz’ equivariant R-group $R^{spin}_{n}(X)^{G}$ depends only on the fundamental groupoid functor of the reference space $X$. The proof covers at the same time in a concise and clear way the classical non-equivariant case.
@@ 1 Introduction
$\dots\to\mathrm{R}^{spin}_{n+1}(X)\to\mathrm{Pos}^{spin}_{n}(X)\to\Omega^{spin}_{n}(X)\to\mathrm{R}^{spin}_{n}(X)\to\dots$ (1.1)
of cobordism groups, where $X$ is a CW-complex. This long exact sequence includes the well-known spin cobordism $\Omega_{*}^{spin}(X)$; then the structure bordism group of metrics of positive scalar curvature $\mathrm{Pos}_{*}^{spin}(X)$, whose cycles $(f\colon M\to X,g)$ are defined like those of $\Omega_{*}^{spin}(X)$, but adding the geometric secondary structure given by a positive scalar curvature metric $g$ on the smooth compact manifold $M$ with spin structure; finally, the object of study of this note, $\mathrm{R}^{spin}_{*}(X)$ which intuitively arises as a sort of mapping cone construction of the forgetful functor $\mathrm{Pos}^{spin}_{*}(X)\to\Omega^{spin}_{*}(X)$.
An explicit calculation of this last group has not yet been achieved in any case of interest. Nonetheless, Stolz has proved that, if $X=B\pi_{1}(M)$ is the classifying space of the fundamental group of a closed spin manifold $M$ with $\dim(M)\geq 5$, then $\mathrm{R}^{spin}_{\dim M+1}(X)$ acts freely and transitively on the space of concordance classes of positive scalar curvature metrics on $M$. This action is canonical and hence the space of concordance classes inherits canonically (after the choice of a base point) an abelian group structure and a construction of an explicit model for this is given in [XieYuZeidler, Section 4]. This has been used, for instance, in [PSZ, PSZ2, SZ, XieYuZeidler] to give a lower bound on the rank of this affine group and of the moduli space of concordance classes of metrics of positive scalar curvature, where moduli space refers to the quotient by the action of the diffeomorphism group of $M$.
A fundamental step in [SZ] uses the fact, which can be implicitly deduced from [Stolz], that a 2-connected map between CW-complexes, such as the classifying map $u\colon M\to B\pi_{1}(M)$, induces an isomorphism between corresponding $\mathrm{R}^{spin}_{*}$ groups for $\geq 6$.
There is an obvious equivariant reformulation of this story: instead of working with a space $X$ one can work with the a $G$-cover $\bar{X}$ with its free $G$-action. One can then go equivalently go back and forth between structure on $X$ and $G$-invariant structure on $\bar{X}$ (pulling back from $X$ to $\bar{X}$ and quotienting by $G$ from $\bar{X}$ to $X$). This way, for a space $X$ with fundamental group $G$ and universal cover $\widetilde{X}$, one can replace equivalently the Stolz sequence (1.1) by its equivariant version
@@
where all the cycles are defined as above, by requiring to come with a free and co-compact and (whenever we have metrics) isometric action of $G$, and all maps are required to be $G$-equivariant. This equivalent reformulation is important to get information about the groups in the Stolz sequence via higher index theory, which has been successfully implemented in [PiazzaSchick] and [Zeidler], compare also [XieYu] and the survey [SchickICM]. The proofs of lower bounds on the rank of the affine group of concordance classes of positive scalar curvature are based on these techniques.
The aim of this note is two-fold. The first main contribution is an explicit and concise proof of the fact that (non-equivariantly) the group $R^{spin}_{*}(X)$ depends for $\geq 6$ only on the fundamental group information of $X$. Secondly, we want to analyze the corresponding statement for the case of general proper $G$-actions. Here, we develop basic tools for this generalization and then study it in the equivariant context of a CW-complex endowed with a proper action of a discrete group. Concretely, for we construct an equivariant version of the Stolz exact sequence, compare Proposition 3.3. One contribution is to give a complete proof of exactness, as we are not aware that this is available in the literature. Our main original result is the fact that the equivariant R-groups do only depend on the equivariant 2-type of the space, compare Theorem 4.1. Along the way, we construct an equivariant analog of the space $B\pi_{1}(X)$, namely a “universal space for a given 2-type, compare Section 5.2.
It should be noted that this relies on surgery constructions requiring enough “room”. This is the reason for the dimension restrictions listed above: the manifolds whose positive scalar curvature metrics are controlled have to be spin manifolds of dimension at least $5$, and the group $R^{spin}_{*}(X)^{G}$ can be treated efficiently if $\geq 6$.
@@ A $G$-CW-complex X is a $G$-space together with a $G$-invariant filtration
Let $f\colon X\to Y$ be a $G$-map. Then there exists a G-homotopy $h\colon X\times I\to Y$ such that $h_{0}:=h_{X\times\{0\}}=f$ and $h_{1}:=h_{X\times\{1\}}$ is cellular.
@@
(compare [ttd, (2.6) Proposition]) Let $f\colon B\to C$ be a $\nu$-connected map between $G$-CW-complexes and $A$ another $G$-CW-complex. Write $A,B]^{G}$ for the set of $G$-homotopy classes of $G$-maps from $A$ to $B$. Then
@@
Note that the classical Whitehead theorem is a consequence, using $A=C$ and $A=B$ and identity maps.
@@ 3 The Stolz exact sequence
  • $R^{spin}_{n}(X)^{G}$ is the bordism group of spin $G$-manifolds with boundary (possibly empty) of dimension $n$, together with a $G$-invariant Riemannian metric of positive scalar curvature on the boundary. Bordisms are then manifolds with corners. In particular, $(M,f,g)$ and $(M^{\prime},f^{\prime},g^{\prime})$ are equivalent if there exists a $G$-bordism $(W,F,\bar{g})$, where $W$ is a bordism between $M$ and $M^{\prime}$ and the resulting bordism $\partial_{0}W$ between $\partial M$ and $\partial M^{\prime}$, carries a $G$-invariant metric $\bar{g}$ with positive scalar curvature, so that it is a bordism between $(\partial M,\partial f,g)$ and $(\partial M^{\prime},\partial f^{\prime},g^{\prime})$ in the sense of $Pos^{spin}_{n-1}(X)^{G}$.
@@
Note that for each cycle $f\colon M\to X$ in $\Omega_{n}^{spin}(X)^{G}$ the isotropy of $M$ is restricted to belong to the family $\mathcal{I}(X)$ (because the image under the equivariant map $f$ of a point $x\in M$ fixed by a subgroup $H$ of $G$ must also be fixed by $H$, and is a point in $X$).
@@
$\ldots\ignorespaces\ignorespaces\ignorespaces\ignorespaces$${R^{spin}_{n+1}(X)^{G}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${Pos^{spin}_{n}(X)^{G}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\Omega^{spin}_{n}(X)^{G}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${R^{spin}_{n}(X)^{G}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\textstyle{\ldots}$
@@
This is a rather direct consequence of the definitions and well known to the experts, at least non-equivariantly, stated e.g. in [Stolz, (4,4)] (but without proof). The argument for the $G$-equivariant case is exactly the same as the classical non-equivariant situation. As we are not aware of a treatment available in the journal literature, we follow the referee’s suggestion and give a complete account of the argument here.
First, that the composition of two consecutive maps is zero indeed is a direct consequence of the definitions:
  • Given $W,f,g]\in R^{spin}_{n+1}(X)^{G}$, its image in $\Omega_{n}(X)^{G}$ is represented by $(\partial W,f|_{\partial W})$ which indeed is a boundary, hence represents 0.
  • Given $M,f,g]\in Pos_{n}^{spin}(X)^{G}$, its image in $R^{spin}_{n}(X)^{G}$ is the manifold $M$ with empty boundary. It is bordant in $R^{spin}_{n}(X)^{G}$ to $\emptyset$ and hence represents 0 via the bordism $M\times[0,1]$, where $(M\times\{1\},g)$ is a positive scalar curvature bordism of $\emptyset=\partial M$ to $\emptyset=\partial\emptyset$.
  • Given $M,f]\in\Omega^{spin}_{n}(X)^{G}$, its image in $Pos_{n-1}^{spin}(X)^{G}$ is represented by $\partial M=\emptyset$, hence represents 0.
For the opposite inclusions of the kernels in the image, the constructions are almost as straight-forward:
  • Assume that $M,f]\in\Omega^{spin}_{n}(X)^{G}$ is mapped to 0 in $R^{spin}_{n}(X)^{G}$. This means that there is a null-bordism $(W,F)$. In the case at hand, $\partial W$ has two disjoint parts: on the one hand $(M,f)$ and on the other hand $(M^{\prime},f^{\prime},g^{\prime})$ which is a bordism from $\partial M=\emptyset$ to $\partial\emptyset=\emptyset$ and which is equipped with a metric $g$ of positive scalar curvature. But this means, by definition, that $M,f]=[M^{\prime},f^{\prime}]\in\Omega^{spin}_{n}(X)^{G}$ where $M^{\prime}$ is equipped with a metric $g$ of positive scalar curvature. Hence we have the preimage $M^{\prime},f^{\prime},g]\in Pos^{spin}_{n}(X)^{G}$ of the initial equivariant bordism class $(M,f)$.
  • Assume that $M,f,g]\in Pos^{spin}_{n}(X)^{G}$ is mapped to 0 in $\Omega^{spin}_{n}(X)^{G}$. This means that there is a bordism $(W,F)$ with $\partial(W,F)=(M,f)$. But then $[W,F,g]$ represents a class in $R^{spin}_{n+1}(X)^{G}$ which is mapped to $M,f,g]\in Pos^{spin}_{n}(X)^{G}$.
  • Finally, assume that $W,f,g]\in R^{spin}_{n+1}(X)^{G}$ is mapped to 0 in $Pos_{n}^{spin}(X)^{G}$. This means that there is a second bordism $(W^{\prime},f^{\prime},g^{\prime})$ whose boundary is $(\partial W,f|_{\partial W},g)$. We now consider the closed $G$-manifold $Z:=W\cup_{\partial W}W^{\prime}$ with map $F:=f\cup_{\partial W}f^{\prime}\colon Z\to X$ and the bordism $B:=Z\times[0,1]$ with map $F\circ pr_{Z}\colon B\to X$. The boundary of $B$ consists of three parts: the internal boundary $W^{\prime}$ between $\partial W$ and $\emptyset=\partial Z$, equipped with the metric $g^{\prime}$ of positive scalar curvature extending $g$. The other two boundary parts are $W$ and $Z$, and by definition $(B,F\circ pr_{Z},g^{\prime})$ is a bordism in $R^{spin}_{n}(X)^{G}$ between $(W,f,g)$ and $(Z,f,\emptyset)$ and $Z,f]\in\Omega^{spin}_{n}(X)^{G}$ represents a preimage of $W,f,g]\in R^{spin}_{n}(X)^{G}$.
@@
    It follows that induction gives an isomorphism (already on the level of cycles and relations)
@@
[…]
This again follows from covering theory: if we fix $x\in X$ with lift $\tilde{x}\in\widetilde{X}$ and a subgroup $H$ of $G$ fixing $x$ we have a canonical split $H\to\widetilde{G}$ sending $h\in H$ to the unique pair $(\alpha\colon\tilde{X}\to\widetilde{X},h)$ such that $\alpha(\tilde{x})=\tilde{x}$.
@@
The action of $\pi_{1}(X)\subset\widetilde{G}$ on $\tilde{X}$ is free. Consequently, for every $\widetilde{G}-$map $f\colon\bar{M}\to\tilde{X}$ the action of $\pi_{1}(X)$ on $\bar{M}$ is free. We can therefore quotient out this action and obtain a cycle $M:=\bar{M}/\pi_{1}(X)\to X$ with residual action of $G:=\widetilde{G}/\pi_{1}(X)$. Vice versa, given a $G$-map $f\colon M\to X$ we can pull back $\tilde{X}\to X$ along $f$ and obtain a $\pi_{1}(X)$-covering $\tilde{M}\to M$ with map $\tilde{f}\colon\tilde{M}\to\tilde{X}$. Because $f$ is a $G$-map, it is straightforward to pull back also the action of $\widetilde{G}$ to an action on $\tilde{M}$ which covers the action of $G$ on $M$ and such that $\tilde{f}$ is a $\widetilde{G}-$map (indeed, the $\widetilde{G}-$action on $\tilde{M}$ is defined mapping a point $(m,\tilde{x})\in\tilde{M}\subset M\times\tilde{X}$ by $(\alpha,g)\in\widetilde{G}$ to $)(gm,\alpha(\tilde{x}$))).
@@
For the paper at hand, the transition to the universal covering as in Proposition 3.6 is not really relevant. However, in other situations this point of view is really fruitful. To our knowledge, essentially all information about non-triviality of the groups $R^{spin}_{n}(X)$ and $Pos_{n}^{spin}(X)$ use higher index theory of the Dirac operator. A particularly transparent way to do this is to use the equivariant Dirac operator on the universal covering. This essentially means that one applies the isomorphism the groups “downstairs” for $X$ and the $\pi_{1}(X)$-equivariant groups “upstairs” for the universal covering $\tilde{X}$, as we have already mentioned in the introduction.
It is pretty obvious that the same observation does hold for non-free actions: even if we want to understand symmetric metrics of positive scalar curvature on a compact manifold $M$ with smooth action of a finite group $\Gamma$, it will be very useful to pass to the covering space which also sees the symmetries induced by the non-trivial fundamental group. This strategy has already been implemented in [XieYu, Section 5] and [XieYuZeidler]. The first paper constructs the map from the equivariant Stolz sequence to analysis (in form of the Higson-Roe exact sequence for the K-theory of the Roe $C^{*}$-algebra) in the sense of [PiazzaSchick]. The second paper then uses this to explicitly distinguish many bordism classes of equivariant positive scalar curvature metrics.
3.1 Refinements beyond bordism of positive scalar curvature metrics
The Stolz positive scalar curvature exact sequence of Proposition 3.3 gives important information about the existence and classification of metrics of positive scalar curvature. However, by the very definition this information is about bordism classes and we are of course also interested in a fixed given manifold $M$.
Non-equivariantly, the situation here is very satisfactory: for a given connected closed spin manifold $M$ with $\dim(M)\geq 5$, if we use $X=B\pi_{1}M$ then $M$ itself admits a metric with positive scalar curvature if (and only if) the image of $u\colon M\to B\pi_{1}(M)]$ in $R_{n}^{spin}(B\pi_{1}(M))$ vanishes. Moreover, in this case $R_{n+1}^{spin}(B\pi_{1}(M))$ acts freely and transitively on the concordance classes of metrics of positive scalar curvature on $M$.
Of course, it would be very desirable to extend such results to the equivariant case. Unfortunately, the proof in the non-equivariant case uses as an important tool “handle cancellation” for Morse functions. It is known that equivariantly such cancellation is not even true. Therefore, a general treatment of the equivariant case seems not in reach at the moment. Under very special conditions on the action, positive existence results can be obtained. The strongest results in this direction we are aware of are obtained in [Hanke]. There, also the difficulties are discussed when one attempts to obtain more general results. In the paper at hand, we focus on the bordism context and do not attempt to contribute to obstruction and classification results for G-invariant metrics of positive scalar curvature on a fixed manifold $M$.